range of function :- x^2+4x
7/x^2+x+1

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
f(x)=x^2+4x+ 7/x^2+x+1
f'(x)=(2x+4)(x^2+x+1)-(2x+1)(x^2+4x+ 7)/(x^2+x+1)^2
f'(x)=0 , gives x=-2
f(-2)=4-8+7/(4-2+1)= 3/3 =1
f(0)=7
range=[1,7]